For all questions in this assignment, consider the 2016 Income Statement and Balance Sheet from SmallCo below.

part 1)

Suppose SmallCo made a large purchase of capital equipment for its distribution centers on Jan 1st 2017 using cash.

How would this purchase affect the income statement and balance sheet for 2017? (Please select all the correct answers)

A) Property, plant, equipment decreases

B) Depreciation increases

C) Property, plant, equipment increases

D) Depreciation decreases

E) None of the above

Part 2)

SmallCo hopes to roll out an inventory reduction program in 2017. The company expects to reduce its inventories by 800 dollars. There has been some concerns that this may affect sales negatively.

If sales were to reduce by 2500 dollars while inventories are reduced by 800 dollars, how would working capital turnover be affected? Assume all other numbers are the same in 2017 as in the 2016 reports. (Please select the one best answer)

A) Working capital turnover will increase by 8%

B) Working capital turnover will increase by 13%

C) Working capital turnover will increase by 17%

D) Working capital turnover will not change

E) None of the above

Part 3) After the inventory reduction project, what would be the cash-to-cash cycle? Assume the COGS, Accounts Receivable, and Accounts Payable in 2017 remain the same as 2016. Answer in full days.

Part 4)

Which of the following statements describes supply chain actions that definitely reduce the cash-to-cash cycle? (Only one correct answer)

A) Reducing the production time in a Make-To-Order setting and receiving payments from customers at the time of order instead of at the time of delivery

B) Increasing inventory levels but delivering faster to customers

C) Increasing inventory levels but demanding faster payments from customers

D) Reducing the production time in a Make-To-Order setting and paying suppliers faster

E) None of the Above
Answer #1

1. Option C. Property, plant, equipment increases

2. Current NWC = Currnt Assets - Current Liabilities =13712 - 6844 = $6868

Current NWC Turnover = Net Sales / NWC = 58643 / 6868 = 8.54 Times

New NWC = Currnt Assets - Current Liabilities =12912 - 6844 = $6068

New NWC Turnover = Net Sales / NWC = 56143 / 6068 = 9.25 Times

Change in NWC Turnover = 9.25 / 8.54 = 8% (Rounded Off) Option A

3. Cash to Cash Cycle

Accounts Receivable Days = (Accounts Receivable / Sales) * 365 = 4026 * 365 / 56143 = 26 Days

Inventory Days = (Inventory / COGS) * 365 = 3965 * 365 / 35243 = 41 Days

Payable Days = (Accounts Payable / COGS) * 365 = 4920 / 35243 = 51 Days

Cash to Cash Cycle = Accounts Receivable Days + Inventory Days - Payable Days

Cash to Cash Cycle = 26 + 41 - 51

Cash to Cash Cycle = 16 Days

4. Option A. Reducing the production time in a Make-To-Order setting and receiving payments from customers at the time of order instead of at the time of delivery